A retrofitted space-heating system is being considered for a small office building. The system can be purchased and installed for $135,000, and it will save an estimated 315,650 kilowatt-hours (kWh) of electric power each year over a eight-year period. A kilowatt-hour of electricity costs $0.15, and the company uses a MARR of 15% per year in its economic evaluations of refurbished systems. The market value of the system will be $10,000 at the end of eight years, and additional annual operating and maintenance expenses are negligible. What is the conventional B/C ratio of this system?

1.41

1.55

1.61

1.68

1.72
Answer #1

Present value of cash inflow / Ini = Present value of saving + Present value of salvage value

= (315650*0.15)*CPVF15%, 8 + $10000*PVF15%, 8

= $47347.50*4.48732 + 10000*0.32690 = $215732.38

Conventional B/C ratio = Present value of cash inflow / Initial Investment

= $215732.38 / $135000

= 1.61